OpenStudy (anonymous):

Shilo is trying to compute the value of 27 + 55 + 13. She adds 27 + 13 to get 40, and then adds on 55. Which of the following properties is Shilo using? commutative distributive associative identity

OpenStudy (anonymous):

Commutative and Associative

OpenStudy (anonymous):

it can only be one

OpenStudy (anonymous):

Commutative - means that you can swap numbers and still get the same thing so: a+b = b+a Distributive - means that when multiplying through a parentheses you must distribute to each term inside the parentheses so: a(b+c) = ab+ac associative - means that it doesn't matter how you group the numbers, it's the same so: (a+b)+c = a + (b+c) identity - simply means that they are equal

OpenStudy (anonymous):

so associative?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Commutative

OpenStudy (anonymous):

Oh yeah. xD

OpenStudy (anonymous):

oh ye makes mmore sence thnks

OpenStudy (anonymous):

lol